OpenStudy (anonymous):

whats the inverse function of the function f(x)= the square root of x+10? algebraically

geerky42 (geerky42):

To find inverse function, you replace f(x) to y, then you switch x and y to each other, then solve for y.

geerky42 (geerky42):

So you have \(f(x) = \sqrt{x+10}\) Replace f(x) to y. \[y = \sqrt{x+10}\] Switch x and y to each other: \[x = \sqrt{y+10}\] Now solve for y. Can you do it from here?